Shift Summary for eve shift, 9/9/99

AG (ag@fnal.gov)
Fri, 10 Sep 1999 00:57:40 -0400 (EDT)

Shift Summary for eve shift, 9/9/99

On Shift: AA, AG

Hours of beam available at all intensities: 3h 30min
Hours of beam available at nominal intensity: 3h 30min
Hours of E832 data (special runs not included): 2h 39min
Hours of special runs (laser scans, muon runs...): 0h 18min

Accounting of lost time:

o Tape drive problems.
o NM4AM problems.

========================================================================

Runs ended this shift:

Run 14486 - 240K events - MUON
Run 14488 - BEAM_832, in progress.

Tapes brought to FCC this shift:
(including tapes for online split)

None.

========================================================================

Accesses during shift:

None.

========================================================================

Overview of shift:

No beam till 2030. When beam returns we took a muon run and then started
taking e832 data. MCR gave us hard time tonight by playing all sorts of
games with magnets.

========================================================================

Description of Shift:

run____spill______time_______comment

------------------------------------------------------------------------

_________________1600______Called MCR to setup muon run configuration.
___________________________They promise beam @ 1630. Trying to initialize
___________________________muon run - tapes come up red. Fresh started...
___________________________Tapes red again - Sydney is working on it.

_________________1700______Called MCR, but the estimate now is 2 hours.

_________________1800______Finally can initialize tapes, but we don't know
___________________________what fixed it. Start COSMICS run.

_________________1930______New crew from MCR called to tell us that they
___________________________set magnets for hadrons, but they suspected
___________________________that we wanted muons. Beam estimate is 1945.
___________________________We clarified that muons were really what we
___________________________wanted and they promised to change magnet
___________________________settings back.
_________________2000______Start muon run... It's aborted. Fresh
___________________________started, initialization problems, SCSI bus on
_________________2035______ktev3 is hunging. We had to cycle power on
___________________________ktev3 tape drives. Meanwhile beam is here.

14486____________2045______Finally started muon run, but trigger rate is
___________________________very low. Apparently, NM2S3 is set to 0 but
___________________________reads ~350. Called MCR - they fixed it - rate
___________________________is back to normal.

14487____________2100______Started BEAM_832 run, but Rick realized that
___________________________E/p plot for Ke3's was screwed up. Turns out
___________________________L3 ptkick was set 0.254 GeV ??

_________________2110______Stop this run, ramp down NM4AM and then ramp it
___________________________up again.

14488____________2121______Start another BEAM_832 run. First 4
___________________________spills got pipe-line errors, everything seems to
___________________________be OK after that.

========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift